# Enhancing Financial Time-Series Forecasting with Retrieval-Augmented Large Language Models

### v2

@@ 3.2. Sequence Serialization
Query Sequence Serialization We select query sequences where the query date is at least one year after the start date of the corresponding split in the dataset. This ensures that each query has enough candidate sequences for retrieval. For example, in the ACL18 dataset, which spans from 2014-06-02 to 2015-12-31, the query sequences are defined as those from 2015-06-03 to 2015-12-31.
Each query sequence contains the adjusted close price from the previous five trading days along with basic stock information. The indicator is represented as a five-day sequence, with a one-day sliding window applied across trading days. For example, when inquerying about stock MO on 2015-06-02, the query sequences contains stock name, query date, last five trading dates and their corresponding adjusted close prices. The serialized sequence is shown below:
@@
Candidate Sequence Serialization As new information becomes available, the candidate pool dynamically incorporates sequences from the most recent trading days. For instance, when the query date is 2015-06-03, the candidate pool includes sequences from 2014-06-02 to 2015-06-02. When the query date advances to 2015-06-04, the sequence for 2015-06-03 is added to the candidate pool without the need for additional training. This demonstrates the scalability of our method.
For each candidate stock on a specific date, we select 6 basic price indicators and 20 groups of financial indicators. Each candidate sequence includes a typical financial indicator along with basic stock information. As the same as query serialization, the indicator is represented as a five-day sequence, with a one-day sliding window applied across trading days. For example, a candidate represent stock MO on date 2014-07-02, with highest price as its indicator. The sequence includes basic information containing stock name, date, movement, last five trading dates and their corresponding highest prices. The serialized sequence is shown below:
@@
3.3. Retriever Training
To identify valuable information from vast amounts of sequences and uncover semantic relationships between them, we introduce a novel retrieval mechanism featuring three key innovations: a fine-tuned 1B parameter LLM (StockLLM) as the backbone, a candidate selection method leveraging LLM feedback to identify beneficial sequences and a training objective that maximizes similarity between queries and historically significant sequences, enabling the trained retriever to uncover meaningful patterns while minimizing noise in complex financial data. While inspired by (Zhang et al., 2023), our approach is uniquely tailored to financial time-series data, addressing the challenges of relevance assessment and alignment in a domain where traditional retrieval methods often fall short. This section details the training process of our framework, including the training of the LLM backbone, candidate scoring and selection, and the training of the retriever, which contains the training objective and knowledge distillation.
The LLM Backbone. To activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA 3.2-1B-Instruct) (Grattafiori et al., 2024) using the LoRA technique for efficient low-rank adaptation, resulting in StockLLM. By intentionally using a smaller backbone model, we establish a more challenging experimental setup, ensuring that performance improvements are attributable to FinSeer’s retrieval capabilities rather than the LLM’s capacity. The fine-tuning process is implemented using the LlamaFactory framework (Zheng et al., 2024), with the following configuration: a learning rate of 5e-5, a cosine scheduler, gradient accumulation over 8 steps, mixed-precision (fp16) training, and 5 epochs of training with regular evaluation to log metrics and select the best-performing checkpoint. This setup demonstrates the robustness of our framework in extracting meaningful insights even under constrained model size, highlighting the effectiveness of FinSeer in enhancing financial forecasting tasks.
Candidate Scoring. To determine whether a candidate sequence assists in predicting the movement of the query, we use LLM feedback to score each candidate. Specifically, for a given query $q$, we integrate the query sequence and each candidate sequence $c_{i}$ from the candidate pool as concurrent inputs to the LLM. The LLM outputs logits, which are unnormalized scores representing the model’s confidence for each possible class (e.g., ”rise” or ”fall”). These logits are transformed into probabilities $P(c)$ using the softmax function:
(4) $P(c)=\frac{e^{z_{c}}}{\sum_{j}e^{z_{j}}},$
where $z_{c}$ is the logit for the correct class (e.g., ”rise” if the true movement is upward) and $z_{j}$ represents the logits for all possible classes. The resulting probability $P(c)$ serves as the score for the candidate $c_{i}$ with respect to the query $q$.
Candidate Selection. We rank the candidate sequences in descending order based on their scores $P(c)$. The top-1 sequence is selected as a positive candidate, while the bottom 15 sequences are chosen as negative candidates. The sets of selected positive and negative sequences are denoted as $\mathbb{C_{P}}$ and $\mathbb{C_{N}}$, respectively.
@@ 3.3. Retriever Training
Knowledge Distillation. To leverage the scoring derived from the LLM, we employ knowledge distillation, which transfers knowledge from the teacher model (LLM) to the student model (retriever) by mimicking the teacher’s output distribution. This approach effectively captures nuanced patterns and predictions from the LLM. Specifically, we minimize the Kullback-Leibler (KL) divergence between the candidate distributions computed using the LLM’s rewards and those predicted by the embedding model. For each query $q$ and its candidate list {$\{\mathbb{C_{P}},\mathbb{C_{N}}\}$, we derive the LLM’s rewards for the candidates, denoted as $\{P(c_{i}),i=1,...,n\}$. To make these rewards suitable for distillation, we normalize them using a softmax function with temperature $\alpha$:
(6) $w_{i}=\text{softmax}_{R}\left(\frac{P(c_{i})}{\alpha}\right).$
@@
(7) $\min\sum_{c}-w_{i}\times\log\left(\frac{\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{c_{i}}\right\rangle/\tau\right)}{\sum_{{c}^{\prime}\in\mathbb{C}}\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{{c}^{\prime}}\right\rangle/\tau\right)}\right),$
@@
3.4. Inference
The inference process of our RAG framework, built on FinSeer, consists of three key steps to enhance stock movement prediction. First, FinSeer retrieves relevant sequences from the candidate pool, leveraging its ability to identify historically significant patterns. Second, these retrieved sequences are incorporated into StockLLM’s input context, enriching the model’s understanding of temporal dynamics. Finally, the enriched context is combined with the original query to generate forecasts. By integrating salient historical patterns with current data, this approach bridges the gap between past trends and present conditions, significantly improving prediction accuracy and robustness in real-world financial scenarios.
4. Experiment
4.1. Experimental Settings
Datasets. We evaluate the effectiveness of our RAG framework on the test sets of the three datasets described in Table 2, with ACL18 containing 2,876 query sequences, BIGDATA22 containing 2,868 queries, and STOCK23 containing 4,128 queries. These thousand-scale queries help mitigate random bias, ensuring a robust and reliable evaluation of model performance. To ensure a comprehensive evaluation, for each query sequence, we include all sequences containing financial indicators across all stocks in the test set (not limited to the same stock), with data available up to the query date, as potential candidates. No additional restrictions are imposed, enabling a robust assessment of the models’ performance in real-world financial forecasting scenarios.
Baselines. To evaluate the efficiency of our RAG framework, we choose bare StockLLM and random retrieval as our baselines. To evaluate our retriever FinSeer, we select five competitive RAG models from the top of the MTEB English Leaderboard as baselines: (1) Instructor-large (Su et al., 2023): A 335M instruction-finetuned text embedder that encodes sequences into 768-dimensional tensors. (2) UAE-large-v1 (Li and Li, 2023): A 335M ANGLE-optimized text embedding model that encodes sequences into 1024-dimensional tensors. (3) BGE-large-en-v1.5 (Xiao et al., 2023): A 335M general embedder pre-trained with RetroMAE (Xiao et al., 2022), encoding sequences into 1024-dimensional tensors. (4) LLM-Embedder (Zhang et al., 2023): A 109M embedder fine-tuned from BGE-large-en-v1.5, designed as a unified embedding model to support diverse retrieval augmentation needs for LLMs. It encodes sequences into 768-dimensional tensors. (5) E5-mistral-7b-instruct (Wang et al., 2023): A 7111M embedder initialized from Mistral-7B-v0.1 (Jiang et al., 2023) and fine-tuned on multilingual datasets, encoding sequences into 4096-di-mensional tensors.
@@ 4.1. Experimental Settings
4.2. Main Results
Table 3 summarizes the performance of stock movement prediction across the three datasets.
Table 3. Results of stock movement predictions using LLMs and retrieval models. The asterisk (*) indicates the LLM employed while using retrieval models.
@@ 4.2. Main Results
(+ StockLLM) ACC MCC ACC MCC ACC MCC
w/o Retrieval 0.496 -0.007 0.497 -0.008 0.509 0.021
Random Retrieve 0.481 -0.036 0.521 0.045 0.496 -0.004
DTW 0.518 0.045 0.505 0.021 0.492 -0.007
Instructor 0.496 -0.008 0.497 -0.006 0.505 0.010
UAE 0.482 -0.036 0.495 -0.008 0.494 -0.009
E5 0.510 0.028 0.503 0.008 0.499 -0.001
BGE 0.489 -0.019 0.487 -0.026 0.488 -0.014
LLM Embedder 0.499 0.000 0.459 -0.083 0.503 0.007
FinSeer 0.521 0.042 0.538 0.079 0.542 0.085
We analyze our framework and retriever separately to evaluate their contributions. First, our RAG framework outperforms both bare StockLLM and random retrieval across all datasets, demonstrating its effectiveness in enhancing stock movement prediction. Second, among all evaluated retrieval methods, FinSeer consistently achieves the best performance when integrated with StockLLM, showcasing its ability to retrieve valuable sequences that improve forecasting accuracy. Notably, FinSeer is the only retriever that achieves stable improvements over bare StockLLM on all datasets, while other methods exhibit inconsistent results, sometimes enhancing predictions and other times degrading them. Furthermore, our experiments reveal that incorporating a retrieval process does not always guarantee performance improvements, as retrieval errors can mislead the LLM and harm predictions. This underscores the critical need for retrieval models that align closely with the LLM’s objectives, a challenge that FinSeer effectively addresses through its tailored design.
Figure 2. Indicator occurrences of different RAG models on ACL18 dataset.
The minimal performance difference between the instruction-finetuned retriever (Instructor) and the no-retrieval baselines highlights the inherent challenge of time-series retrieval. Unlike text retrieval, time-series retrieval does not solely rely on task understanding, as candidate sequences often appear similar, making it difficult to distinguish valuable ones. The angle-optimized retriever (UAE) shows inconsistent performance across datasets, as its ability to differentiate similarity between query and candidate sequences does not fully align with the LLM’s perception of importance. Similarly, LLM Embedder, our backbone model trained with LLM feedback, struggles to generalize to time-series retrieval tasks, further underscoring the complexity of the problem.
The best-performing baseline is E5, which achieves competitive results despite its significantly larger model size and embedding dimensions (five times larger than FinSeer). However, FinSeer outperforms E5 on all datasets. This demonstrates that FinSeer successfully learns the preferences of LLMs and effectively enhances their time-series forecasting ability through retrieval, even under challenging conditions.
In summary, while RAG methods do not guarantee performance improvements due to the risk of retrieval errors, FinSeer stands out by consistently retrieving high-quality sequences that align with the LLM’s forecasting objectives, leading to superior results compared to other retrievers.
4.3. Ablation Study
Indicator Occurrences In this part, we analyze what indicators all RAG models retrieve. We calculate indicator occurrences on ACL18 dataset and the results are shown in Figure 2.

### v3

@@ 3.1.1 Sequence Serialization
The following are two examples of a query and a candidate sequence. When inquerying about stock MO on 2015-06-02, the query sequence contains the stock identifier (MO), query date (2015-06-02), and the adjusted close prices of last five trading dates. The serialized sequence is shown below:
@@
A potential candidate in the candidate pool represent stock MO on date 2014-07-02, with the highest price as its indicator. The sequence includes the stock ticker (MO), price movement direction (freeze) on a specific trading date (2014-07-02), a key financial indicator (the highest price), and corresponding indicator values calculated from the preceding five consecutive trading days. The serialized sequence is shown below:
@@
3.2 Retriever Training
We then train FinSeer to effectively distinguish high-value sequences (positives) from irrelevant sequences (negatives) by learning a retrieval embedding space where relevant candidates are closely aligned with their corresponding queries. To achieve this, we score and select positive and negative candidates as the retriever training corpus (in Candidate Scoring and Selection), then specify the training objective (in Training Objective) and conduct knowledge distillation (in Knowledge Distillation).
Candidate Scoring and Selection. To determine whether a candidate sequence assists in predicting the movement of the query, we use LLM feedback to score each candidate. The details of our LLM backbone, StockLLM-1B-Instruct (hereafter referred to as StockLLM), are shown in Appendix F.1. Specifically, for a given query $q$, we integrate the query sequence and each candidate sequence $c_{i}$ from the candidate pool as concurrent inputs to the StockLLM. Then StockLLM outputs logits, which are unnormalized scores representing the model’s confidence for each possible class (e.g., rise or fall). These logits are transformed into probabilities $P(c)$ using the softmax function:
$P(c)=\frac{e^{z_{c}}}{\sum_{j}e^{z_{j}}},$ (1)
where $z_{c}$ is the logit for the correct class (e.g., rise if the true movement is upward) and $z_{j}$ represents the logits for all possible classes. The resulting probability $P(c)$ serves as the score for the candidate $c_{i}$ with respect to the query $q$.
We rank the candidate sequences in descending order based on their scores $P(c)$. The top-1 sequence is selected as a positive candidate, while the bottom 15 sequences are chosen as negative candidates. The sets of selected positive and negative sequences are denoted as $\mathbb{C_{P}}$ and $\mathbb{C_{N}}$, respectively.
@@ 3.2 Retriever Training
Knowledge Distillation. To leverage the scoring derived from StockLLM, we employ knowledge distillation, which transfers knowledge from the teacher model (StockLLM) to the student model (retriever) by mimicking the teacher’s output distribution. This approach effectively captures nuanced patterns and predictions from StockLLM. Specifically, we minimize the Kullback-Leibler (KL) divergence between the candidate distributions computed using the LLM’s rewards and those predicted by the embedding model. For each query $q$ and its candidate list {$\{\mathbb{C_{P}},\mathbb{C_{N}}\}$, we derive StockLLM’s rewards for the candidates, denoted as $\{P(c_{i}),i=1,...,n\}$. To make these rewards suitable for distillation, we normalize them using a softmax function with temperature $\alpha$:
$w_{i}=\text{softmax}_{R}\left(\frac{P(c_{i})}{\alpha}\right).$ (3)
@@
$\min\sum_{c}-w_{i}\times\log\left(\frac{\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{c_{i}}\right\rangle/\tau\right)}{\sum_{{c}^{\prime}\in\mathbb{C}}\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{{c}^{\prime}}\right\rangle/\tau\right)}\right),$ (4)
@@
3.3 Inference
During inference, the key innovation of our FinSrag framework lies in how FinSeer’s retrieval directly enhances StockLLM’s forecasting capability. Given a query, FinSeer first identifies the most relevant historical sequences by evaluating both temporal patterns and predictive relationships learned during training, filtering out noisy but numerically similar candidates that typically mislead traditional retrievers. These selected sequences are then structured and injected into StockLLM’s context window. Crucially, unlike standard RAG that simply concatenates retrieved documents, this end-to-end alignment between retrieval and generation is what enables FinSrag to outperform conventional forecasting pipelines where retrieval and prediction models are optimized separately.
4 Experiment
4.1 Experimental Settings
Datasets. We evaluate the effectiveness of our RAG framework on the test sets of the three datasets described in Table 5, with ACL18 containing 2,876 query sequences, BIGDATA22 containing 2,868 queries, and STOCK23 containing 4,128 queries. These thousand-scale queries help mitigate random bias, ensuring a robust and reliable evaluation of model performance.
Candidate Pool Settings. To ensure a comprehensive evaluation, for each query sequence, we include all sequences containing financial indicators across all stocks in the test set (not limited to the same stock), with data available up to the query date, as potential candidates. No additional restrictions are imposed, enabling a robust assessment of the models’ performance in real-world financial forecasting scenarios.
Baselines. To evaluate the efficiency of the FinSrag framework, the bare StockLLM-1B-Instruct without retrieval serves as our baseline to figure out whether the retrieval step enhances StockLLM’s prediction ability. To evaluate our retriever FinSeer, we tested other retrieval methods, including random retrieval, DTW distance, and five competitive retrieving models from the top of the MTEB English Leaderboard as baselines, containing: (1) Instructor-large (Su et al., 2023), a 335M instruction-finetuned text embedder that encodes sequences into 768-dimensional tensors. (2) UAE-large-v1 (Li and Li, 2023), a 335M ANGLE-optimized text embedding model that encodes sequences into 1024-dimensional tensors. (3) E5-mistral-7b-instruct (Wang et al., 2023), a 7111M embedder initialized from Mistral-7B-v0.1 (Jiang et al., 2023) and fine-tuned on multilingual datasets, encoding sequences into 4096-di-mensional tensors. (4) BGE-large-en-v1.5 (Xiao et al., 2023), a 335M general embedder pre-trained with RetroMAE (Xiao et al., 2022), encoding sequences into 1024-dimensional tensors. (5) LLM-Embedder (Zhang et al., 2023), a 109M embedder fine-tuned from BGE-large-en-v1.5, designed as a unified embedding model to support diverse retrieval augmentation needs for LLMs. It encodes sequences into 768-dimensional tensors. More details are shown in Appendix F.2.
@@ 4.1 Experimental Settings
4.2 Main Results
As shown in Table 2, experimental results demonstrate our framework’s effectiveness by outperforming all baseline retrieval methods in assisting LLM-based stock movement prediction.
Table 2: Results of stock movement predictions using StockLLM-1B-Instruct and retrieval models.
@@ 4.2 Main Results
(+ StockLLM-1B-Instruct ) ACC MCC ACC MCC ACC MCC
w/o Retrieval 0.498 -0.006 0.493 -0.017 0.509 0.021
Random Retrieval 0.485 -0.028 0.495 -0.007 0.496 -0.004
DTW 0.516 0.041 0.500 0.010 0.492 -0.007
Instructor (Su et al., 2023) 0.498 -0.005 0.493 -0.015 0.505 0.010
UAE (Li and Li, 2023) 0.486 -0.029 0.493 -0.011 0.494 -0.009
E5 (Wang et al., 2023) 0.510 0.027 0.498 -0.002 0.499 -0.001
BGE (Xiao et al., 2023) 0.492 -0.012 0.501 0.002 0.488 -0.014
LLM Embedder (Zhang et al., 2023) 0.503 0.007 0.459 -0.083 0.503 0.007
FinSeer 0.517 0.035 0.510 0.023 0.542 0.085
First, compared with bare StockLLM and random retrieval, FinSeer demonstrates the critical importance of retrieving truly valuable information. While trained on a comprehensive stock movement prediction corpus, bare StockLLM’s limited input with only the recent five-day adjusted close price results in unstable performance that fluctuates around random guessing levels. Similarly, when randomly retrieved sequences are provided as supposedly relevant context, they introduce instability by arbitrarily confusing or occasionally coincidentally benefiting StockLLM’s decision-making.
Second, our comparison with the five top-ranked retrievers from the MTEB English leaderboard reveals FinSeer’s superiority in time-series retrieval. The negligible performance gap between instruction-finetuned retriever (Instructor) and no-retrieval baselines underscores the fundamental challenges of time-series retrieval. Unlike text retrieval, this task cannot rely solely on task understanding since candidate sequences often exhibit visual similarity while differing in predictive value. Other retrievers demonstrate inconsistent cross-dataset performance because their similarity differentiation fails to align with the LLM’s perception of importance. Even LLM Embedder, our backbone model trained with LLM feedback, shows limited generalization to time-series retrieval, further emphasizing the problem’s complexity. Among those retrievers, FinSeer consistently outperforms other retrievers across all datasets, proving its superior ability to learn LLM preferences and effectively enhance time-series forecasting through retrieval.
Third, our framework’s advantages over distance-based retrieval methods highlight the value of incorporating diverse feature types. While DTW achieves comparable performance to FinSeer during specific market periods (ACL18 [2014-2015] and BIGDATA22 [2019-2020]), its retrieval capability proves insufficient during the volatile 2022-2023 period (STOCK23) marked by significant market surges and fluctuations. During this challenging phase, DTW significantly degrades StockLLM’s performance, while FinSeer maintains its enhancement capability.
4.3 Ablation Study
In this section, we explore two aspects of our framework based on retrieval results: which indicators are most retrieved by all retrievers (in Indicator Occurrences), and whether StockLLM-1B-Instruct makes predictions by analyzing candidates or just based on the candidates’ movements (in Appendix G.1). We also visualize indicator sequence embeddings in Appendix G.2. Moreover, we explore the performance of these retrieval methods with a larger size of StockLLM (in Appendix G.3).